helium or nitogen travel faster
1
Expert's answer
2020-09-10T14:39:51-0400

Explanations & Calculations


First check

  • By "\small PV = nRT" & "\small PV = \frac{1}{3} mN\bar{C^2}" a value for the speed of a given gas can be derived as "\small \bar{C^2} = \large\sqrt{\frac{3RT}{M}}"
  • M = molar mass of the considered gas
  • MHelium < MNitrogen
  • Therefore, it could be seen that at a given temperature, the speed of Helium is greater than that of Nitrogen.


Second check

  • Escape speed at the earth's surface is about 11kmh-1 meaning any substance which possess a speed equal or greater to that value, is capable of leaving the earth.
  • Its seen that the percentage of Helium is very low compared to that of Nitrogen, which means Helium has a average speed near that value than Nitrogen at a given temperature (ex : room temperature).
  • This again explains that Helium has a greater speed while Nitrogen has a comparatively very low value.
  • Nitrogen - 78% while Helium < 0.97% in the atmosphere.